Assign a base peak, molecular ion peak (with m/z) and identify the structure.
100- MS-N-0025 Relative Intensity 10 2030 70 80 90 m /z rse Cha O a OOTW
Answer #1

im The peak ~88 m/z is the molecular, peak. & the base peak it at m/2=43 (or its intensity in max.) So, 88-43 = 45 ha loss